# Generative Uncertainty in Diffusion Models

### v1

@@ 4.1 Detecting Low-Quality Generations
Table 1: Image generation results for 10K filtered samples (out of 12K). Our generative uncertainty outperforms previously proposed uncertainty-based approaches in terms of image quality (AU [De Vita and Belagiannis, 2025], BayesDiff [Kou et al., 2024]), as indicated by higher FID and precision scores, and is competitive with non-uncertainty methods (Realism [Kynkäänniemi et al., 2019], Rarity [Han et al., 2023]). We report mean values along with standard deviations over 3 runs with different random seeds.
ADM (DDIM), ImageNet 128×128 UViT (DPM), ImageNet 256×256
Random $11.31\pm 0.07$ $58.90\pm 0.36$ $70.68\pm 0.38$ $9.46\pm 0.12$ $60.94\pm 0.24$ $73.82\pm 0.33$
BayesDiff $11.20\pm 0.05$ $58.80\pm 0.05$ $70.62\pm 0.32$ $9.16\pm 0.17$ $61.77\pm 0.19$ $73.72\pm 0.38$
AU $11.39\pm 0.05$ $58.82\pm 0.42$ $70.70\pm 0.38$ $9.20\pm 0.12$ $61.80\pm 0.33$ $73.46\pm 0.24$
[…]
Realism $9.76\pm 0.04$ $67.95\pm 0.19$ $66.32\pm 0.40$ $8.24\pm 0.09$ $70.29\pm 0.15$ $69.12\pm 0.32$
Rarity $10.09\pm 0.02$ $64.99\pm 0.16$ $67.73\pm 0.47$ $8.37\pm 0.11$ $67.21\pm 0.10$ $67.76\pm 0.48$
@@
To evaluate whether our newly introduced generative uncertainty can be used to detect low-quality generations, we follow the experimental setup from prior work on uncertainty-based filtering [Kou et al., 2024, De Vita and Belagiannis, 2025]. Specifically, we generate 12K samples using a given diffusion model and compute the uncertainty estimate for each sample. We then select the 10K samples with the lowest uncertainty. If uncertainty reliably reflects the visual quality of generated samples, filtering based on it should yield greater improvements in population-level metrics (such as FID) compared to selecting a random subset of 10K images.
Implementation Details
To ensure a fair comparison with BayesDiff [Kou et al., 2024], we adopt their proposed implementation of the last-layer Laplace approximation. Specifically, we use an Empirical Fisher approximation of the Hessian with a diagonal factorization [Daxberger et al., 2021a]. When computing the posterior predictive distribution (Eq. 3.3), we use $M=5$ Monte Carlo samples. For the semantic feature extractor $c_{\phi}$, we leverage a pretrained CLIP encoder [Radford et al., 2021]. Additional implementation details are provided in Appendix B.
Baselines We first compare our proposed generative uncertainty to existing uncertainty-based approaches for detecting low-quality samples: BayesDiff and the aleatoric uncertainty (AU) approach proposed by De Vita and Belagiannis [2025]. BayesDiff estimates epistemic uncertainty in diffusion models using a last-layer Laplace approximation and tracks this uncertainty throughout the entire sampling process. In contrast, in AU aleatoric uncertainty is computed by measuring the sensitivity of intermediate diffusion scores to random perturbations. Unlike our approach, both methods estimate uncertainty directly in pixel space.
Importantly, we also compare our method against non-uncertainty-based sample-level metrics, such as the realism score [Kynkäänniemi et al., 2019] and the rarity score [Han et al., 2023]. These metrics work by measuring the distance of a generated sample from the data manifold (derived from a reference dataset) in a semantic space spanned by the inception-net features [Szegedy et al., 2016]. Notably, prior work [Kou et al., 2024, De Vita and Belagiannis, 2025] has not considered such comparisons, which we believe are essential for assessing the practical utility of uncertainty-based filtering.
Evaluation Metrics
In addition to the widely used Fréchet Inception Distance (FID) [Heusel et al., 2017] for evaluating the quality of a filtered set of images, we also report precision and recall metrics [Sajjadi et al., 2018, Kynkäänniemi et al., 2019]. To compute these quantities we fit two manifolds in feature space: one for the generated images and another for the reference (training) images. Precision is the proportion of generated images that lie in the reference image manifold while recall is the proportion of reference images that lie in the generated image manifold. Precision measures the quality (or fidelity) of generated samples, whereas recall quantifies their diversity (or coverage over the reference distribution).
Results
We present our main results on the ImageNet dataset in Table 1. We first observe that existing uncertainty-based approaches (BayesDiff and AU) result in little to no improvement in metrics that assess sample quality (FID and precision). In contrast, our generative uncertainty method leads to significant improvements in terms of both FID and precision. For example, on the UViT model [Bao et al., 2023], a subset of images selected based on our uncertainty measure achieves an FID of $7.89$, significantly outperforming both the Random baseline ($9.45$) and existing uncertainty-based methods (BayesDiff $9.16$, AU $9.20$).
Next, in order to qualitatively demonstrate the effectiveness of our approach, we show 25 samples with the highest and lowest generative uncertainty (out of the original 12K samples) according to our method in Figure 2. High-uncertainty samples exhibit numerous artefacts, and in most cases, it is difficult to determine what exactly they depict. Combined with the quantitative results in Table 1, this supports our hypothesis that (Bayesian) generative uncertainty is an effective metric for identifying low-quality samples. Conversely, the lowest-uncertainty samples are of high quality, with most appearing as ‘canonical’ examples of their respective (conditioning) class.
For comparison, in Figure 5 we also depict the 25 ‘worst’ and ‘best’ samples according to the uncertainty estimate from BayesDiff. It is evident that their uncertainty is less informative for sample quality than ours. Moreover, their uncertainty measure appears to be very sensitive to the background pixels. Most images with the highest uncertainty have a ‘cluttered’ background, whereas most images with the lowest uncertainty have a ‘clear’ background. We attribute this issue to the fact that in BayesDiff the uncertainty is computed directly in the pixel space, unlike in our approach where we use the semantic likelihood (Section 3.3) to move away from the (high-dimensional) sample space. To further verify the importance of the semantic likelihood, in Figure 7 we perform an ablation where we compute our generative uncertainty directly in the pixel-space. It is clear that without semantic likelihood, our uncertainty becomes overly sensitive to the background pixels in the same way as in BayesDiff.
@@ Results
Returning to Table 1, we observe that filtering based on our generative uncertainty results in some loss of sample diversity, as evidenced by lower recall scores (e.g., $73.82$ for Random vs. $71.92$ for our method on the UViT model). We attribute this to the fact that, in our main experiment, 12K images are generated such that all 1000 ImageNet classes are represented. [^2] Following Kou et al. [2024], we use class-conditional diffusion models but randomly sample a class for each of the 12K generated samples. Since certain classes produce images with higher uncertainty (see Appendix A.6 for a detailed analysis), filtering based on uncertainty inevitably alters the class distribution among the selected samples. Moreover, the trade-off between improving sample quality (precision) and reducing diversity (recall) has been observed before, see for example the literature on classifier-free guidance [Ho and Salimans, 2022].
Lastly, we compare our proposed method with non-uncertainty-based approaches—a comparison missing in prior literature [Kou et al., 2024, De Vita and Belagiannis, 2025]. For realism, we retain the 10K images with the highest scores, whereas for rarity, we keep those with the lowest scores. As shown in Table 1, our generative uncertainty is the only uncertainty-based method that approaches realism and rarity in terms of FID (e.g., $7.89$ for ours vs. $8.24$ for realism and $8.37$ for rarity on UViT). However, a large gap remains in precision (e.g., $64.14$ for ours vs. $70.29$ for realism and $67.21$ for rarity on UViT). Notably, realism and rarity sacrifice the most sample diversity, as indicated by their lowest recall scores (e.g., $69.12$ for realism and $67.76$ for rarity on UViT).
@@
Furthermore, Table 2 shows that our score can be effectively combined with realism or rarity scores. Specifically, combining our score with realism yields an FID of $7.60$ on UViT, compared to $8.26$ when combining realism and rarity. We attribute higher benefits from ensembling our score to the fact that, while realism and rarity exhibit a strong negative Spearman correlation (-0.85), our uncertainty measure is less correlated with them (-0.27 with realism, 0.38 with rarity), as shown in Figure 10. Taken together, these results indicate that our uncertainty score captures (somewhat) different desirable properties of images compared to realism and rarity.
@@ 5 Related Work
Also related is DECU [Berry et al., 2024], which employs an efficient variant of deep ensembles [Lakshminarayanan et al., 2017] to capture the epistemic uncertainty of conditional diffusion models. However, DECU does not consider using uncertainty to detect poor-quality generations, as its framework provides uncertainty estimates at the level of the conditioning variable, whereas ours estimates uncertainty at the level of initial random noise. Similarly, in Chan et al. [2024] the use of hyper-ensembles is proposed to capture epistemic uncertainty in diffusion models for inverse problems such as super-resolution, but, as in DECU, their approach does not provide uncertainty estimates in unconditional settings or in conditional settings with low-dimensional conditioning (such as class-conditional generation). Moreover, both DECU [Berry et al., 2024] and Chan et al. [2024] require modifying and retraining diffusion model components, whereas our approach operates post-hoc with any pretrained diffusion model via the Laplace approximation [Daxberger et al., 2021a]. A recent approach, PUNC [Franchi et al., 2024], focuses only on text-to-image models. The uncertainty of image generation with respect to text conditioning is measured through the alignment between a caption generated from a generated image and the original prompt used to generate said image.
@@
Bayesian inference in generative models has been explored previously outside the domain of diffusion models. Prominent examples include Saatci and Wilson [2017] where a Bayesian version of a GAN is proposed, showing improvements for semi-supervised learning, and Daxberger and Hernández-Lobato [2019], where a Bayesian VAE [Tran et al., 2023] is shown to provide more informative likelihood estimates for the unsupervised out-of-distribution detection compared to the non-Bayesian counterparts [Nalisnick et al., 2018]. Since diffusion models can be interpreted as neural ODEs [Song et al., 2020b], another relevant work is Ott et al. [2023], which employs a Laplace approximation to quantify uncertainty when solving neural ODEs [Chen et al., 2018]. However, Ott et al. [2023] focuses solely on low-dimensional regression problems.
Non-uncertainty based approaches for filtering out poor generations include the realism [Kynkäänniemi et al., 2019], rarity [Han et al., 2023], and anomaly scores [Hwang et al., 2024]. Our work is the first to establish a connection between these scores and uncertainty-based methods, which we hope will inspire the development of even better sample-level metrics in the future. Additionally, a large body of work focuses on specially designed sample-quality scoring models [Gu et al., 2020, Zhao et al., 2024] or, alternatively, on leveraging large pretrained vision-language models (VLMs) [Zhang et al., 2024] for scoring generated images. However, these approaches require either access to sample-quality labels or rely on (expensive) external VLMs. In contrast, our uncertainty-based method requires neither, making it a more accessible and scalable alternative.
@@ 6 Limitations
While we have demonstrated in Section 4 that semantic likelihood is essential for addressing the over-sensitivity of prior work to background pixels [Kou et al., 2024], our reliance on a pretrained image encoder like CLIP [Radford et al., 2021] limits the applicability of our diffusion uncertainty framework to natural images. Removing the dependence on such encoders would unlock the application our Bayesian framework to other modalities where diffusion models are used, such as molecules [Hoogeboom et al., 2022, Cornet et al., 2024] or text [Gong et al., 2022, Yi et al., 2024]. Exploring whether insights from the literature on uncovering semantic features in diffusion models [Kwon et al., 2022, Luo et al., 2024, Namekata et al., 2024] could help achieve this represents a promising direction for future work.
@@ 7 Conclusion
We introduced generative uncertainty and demonstrated how to estimate it in modern generative models such as diffusion. Our experiments showed the effectiveness of generative uncertainty in filtering out low-quality samples. For future work, it would be interesting to explore broader applications of Bayesian principles in generative modeling beyond detecting poor-quality generations. Promising directions include guiding synthetic data generation, detecting memorized samples, and optimizing diffusion hyperparameters via the marginal likelihood using the Laplace approximation.
@@ Acknowledgements.
This project was generously supported by the Bosch Center for Artificial Intelligence. Eric Nalisnick did not utilize resources from Johns Hopkins University for this project. Stephan Mandt acknowledges funding from the National Science Foundation (NSF) through an NSF CAREER Award IIS-2047418, IIS-2007719, the NSF LEAP Center, the IARPA WRIVA program, and the Hasso Plattner Research Center at UCI.
@@
  • In Appendix A.1, we qualitatively compare our method with BayesDiff [Kou et al., 2024].
  • In Appendix A.2, we perform ablations on our semantic likelihood (Section 3.3).
  • In Appendix A.3, we demonstrate how to use our generative uncertainty for pixel-wise uncertainty.
@@
  • In Appendix A.4, we show that diffusion’s own likelihood is not useful for filtering out poor samples.
@@
  • In Appendix A.5, we further compare our generative uncertainty to realism [Kynkäänniemi et al., 2019] and rarity [Han et al., 2023] scores.
@@
  • In Appendix A.6, we investigate the drop in sample diversity by looking at the average generative uncertainty per conditioning class.
@@
  • In Appendix A.7, we apply our generative uncertainty to detect low-quality samples in a latent flow matching model [Dao et al., 2023].
  • In Appendix B, we provide implementation and experimental details.
Appendix A Additional Results
A.1 Qualitative Comparison with BayesDiff

### v2

@@ 4.2 Detecting Low-Quality Generations
Figure 3: Image generation results for $n\in\{6\textrm{K},7\textrm{K},\ldots,11\textrm{K}\}$ filtered samples (out of 12K) for UViT diffusion model [Bao et al., 2023]. Our generative uncertainty outperforms previously proposed uncertainty-based approaches (AU [De Vita and Belagiannis, 2025], BayesDiff [Kou et al., 2024]) in terms of image quality, as indicated by higher FID (left) and precision (middle) scores. We report mean values along with standard deviations over 5 runs with different random seeds.
To demonstrate that our proposed generative uncertainty is effective for detecting low-quality generations also on high-dimensional data such as natural images, we follow the experimental setup from prior work on uncertainty-based filtering [Kou et al., 2024, De Vita and Belagiannis, 2025]. Specifically, we generate 12K samples using a given diffusion model and compute the uncertainty estimate for each sample. We then select $n\in\{6\textrm{K},7\textrm{K},\ldots,11\textrm{K}\}$ samples with the lowest uncertainty. If uncertainty reliably reflects the visual quality of generated samples, filtering based on it should yield greater improvements in population-level metrics (such as FID) compared to selecting a random subset of $n$ images.
Implementation Details
To ensure a fair comparison with BayesDiff [Kou et al., 2024], we adopt their proposed implementation of the last-layer Laplace approximation. Specifically, we use an Empirical Fisher approximation of the Hessian with a diagonal factorization [Daxberger et al., 2021a]. When computing the posterior predictive distribution (Eq. 3.3), we use $M=5$ Monte Carlo samples. For the semantic feature extractor $c_{\phi}$, we leverage a pretrained CLIP encoder [Radford et al., 2021]. We set the observation noise to $\sigma^{2}=0.001$ (Eq. 7). Additional implementation details are provided in Appendix D.
Baselines We compare our proposed generative uncertainty to existing uncertainty-based approaches for detecting low-quality samples: BayesDiff and the aleatoric uncertainty (AU) approach proposed by De Vita and Belagiannis [2025]. BayesDiff estimates epistemic uncertainty in diffusion models using a last-layer Laplace approximation and tracks this uncertainty throughout the entire sampling process. In contrast, in AU, uncertainty is computed by measuring the sensitivity of intermediate diffusion scores to random perturbations. Unlike our approach, both methods estimate uncertainty directly in pixel space.
Evaluation Metrics
In addition to the widely used Fréchet Inception Distance (FID) [Heusel et al., 2017] for evaluating the quality of a filtered set of images, we also report precision and recall metrics [Sajjadi et al., 2018, Kynkäänniemi et al., 2019]. To compute these quantities we fit two manifolds in feature space: one for the generated images and another for the reference (training) images. Precision is the proportion of generated images that lie in the reference image manifold, while recall is the proportion of reference images that lie in the generated image manifold. Precision measures the quality (or fidelity) of generated samples, whereas recall quantifies their diversity (or coverage over the reference distribution).
Results
We present our main results on the ImageNet dataset in Figure 3 for UViT model [Bao et al., 2023] and in Figure 7 for ADM model [Dhariwal and Nichol, 2021]. We first observe that existing uncertainty-based approaches (BayesDiff and AU) result in little to no improvement in metrics that assess sample quality (FID and precision). In contrast, our generative uncertainty method leads to significant improvements in terms of both FID and precision. For example, on the UViT model, a subset of $n=10\textrm{K}$ images selected based on our uncertainty measure achieves an FID of $7.89$, significantly outperforming both the Random baseline ($9.45$) and existing uncertainty-based methods (BayesDiff $9.16$, AU $9.20$).
Next, in order to qualitatively demonstrate the effectiveness of our approach, we show 25 samples with the highest and lowest generative uncertainty (out of the original 12K samples) according to our method in Figure 4. High-uncertainty samples exhibit numerous artefacts, and in most cases, it is difficult to determine what exactly they depict. Combined with the quantitative results in Figures 3&7, this supports our hypothesis that (Bayesian) generative uncertainty is an effective metric for identifying low-quality samples. Conversely, the lowest-uncertainty samples are of high quality, with most appearing as ‘canonical’ examples of their respective (conditioning) class.
For comparison, in Figure 8 we also depict the 25 ‘worst’ and ‘best’ samples according to the uncertainty estimate from BayesDiff. It is evident that their uncertainty is less informative for sample quality than ours. Moreover, their uncertainty measure appears to be very sensitive to the background pixels. Most images with the highest uncertainty have a ‘cluttered’ background, whereas most images with the lowest uncertainty have a ‘clear’ background. We attribute this issue to the fact that in BayesDiff the uncertainty is computed directly in the pixel space, unlike in our approach where we use the semantic likelihood (Section 3.3) to move away from the (high-dimensional) sample space. To further verify the importance of the semantic likelihood, we perform an ablation where we compute the generative uncertainty directly in the pixel-space (see lines). As seen in Figures 3&7, this leads to worse FID/precision numbers in most cases compared to using the proposed semantic likelihood. Moreover, based on samples in Figure 10, it is clear that without semantic likelihood, our uncertainty becomes overly sensitive to the background pixels in the same way as in BayesDiff.
@@ Results
Lastly, we observe that filtering based on our generative uncertainty results in some loss of sample diversity, as evidenced by lower recall scores (see right plots in Figures 3&7). We attribute this to the fact that, in our main experiment, 12K images are generated such that all 1000 ImageNet classes are represented. [^2] Following Kou et al. [2024], we use class-conditional diffusion models but randomly sample a class for each of the 12K generated samples. Since certain classes produce images with higher uncertainty (see Appendix C.6 for a detailed analysis), filtering based on uncertainty inevitably alters the class distribution among the selected samples. Moreover, the trade-off between improving sample quality (precision) and reducing diversity (recall) has been observed before, see for example the literature on classifier(-free) guidance [Ho and Salimans, 2022].
Comparison with realism and rarity scores
Lastly, we compare our proposed method with non-uncertainty-based approaches, such as the realism score [Kynkäänniemi et al., 2019] and the rarity score [Han et al., 2023]. These metrics work by measuring the distance of a generated sample from the data manifold (derived from a reference dataset) in a semantic space spanned by the inception-net features [Szegedy et al., 2016]. Notably, prior work [Kou et al., 2024, De Vita and Belagiannis, 2025] has not considered such comparisons, which we believe are essential for assessing the practical utility of uncertainty-based filtering.
For realism, we retain the $n$ images with the highest scores, whereas for rarity, we keep those with the lowest scores. As shown in Figures 13&14, our generative uncertainty performs on par with realism and rarity in terms of FID. However, compared to our method, realism and rarity scores result in sharper precision-recall trade-offs—yielding larger precision gains at the expense of a greater drop in recall.
@@
Furthermore, Table 1 shows that our score can be effectively combined with realism or rarity scores. Specifically for $n=10\textrm{K}$, combining our score with realism yields an FID of $7.60$ on UViT, compared to $8.26$ when combining realism and rarity. We attribute higher benefits from ensembling our score to the fact that, while realism and rarity exhibit a strong negative Spearman correlation (-0.85), our uncertainty measure is less correlated with them (-0.27 with realism, 0.38 with rarity), as shown in Figure 15. Taken together, these results indicate that our uncertainty score captures (somewhat) different desirable properties of images compared to realism and rarity.
@@ 5 Related Work
Also related is DECU [Berry et al., 2024], which employs an efficient variant of deep ensembles [Lakshminarayanan et al., 2017] to capture the epistemic uncertainty of conditional diffusion models. However, DECU does not consider using uncertainty to detect poor-quality generations, as its framework provides uncertainty estimates at the level of the conditioning variable, whereas ours estimates uncertainty at the level of initial random noise. Similarly, in Chan et al. [2024] the use of hyper-ensembles is proposed to capture epistemic uncertainty in diffusion models for inverse problems such as super-resolution, but, as in DECU, their approach does not provide uncertainty estimates in unconditional settings or in conditional settings with low-dimensional conditioning (such as class-conditional generation). Moreover, both DECU [Berry et al., 2024] and Chan et al. [2024] require modifying and retraining diffusion model components, whereas our approach operates post-hoc with any pretrained diffusion model via the Laplace approximation [Daxberger et al., 2021a]. A recent approach, PUNC [Franchi et al., 2025], focuses only on text-to-image models. The uncertainty of image generation with respect to text conditioning is measured through the alignment between a caption generated from a generated image and the original prompt used to generate said image.
@@
Bayesian inference in generative models has been explored previously outside the domain of diffusion models. Prominent examples include Saatci and Wilson [2017] where a Bayesian version of a GAN is proposed, showing improvements for semi-supervised learning, and Daxberger and Hernández-Lobato [2019], where a Bayesian VAE [Tran et al., 2023] is shown to provide more informative likelihood estimates for the unsupervised out-of-distribution detection compared to the non-Bayesian counterparts [Nalisnick et al., 2019]. Since diffusion models can be interpreted as neural ODEs [Song et al., 2021c], another relevant work is Ott et al. [2023], which employs a Laplace approximation to quantify uncertainty when solving neural ODEs [Chen et al., 2018]. However, Ott et al. [2023] focuses solely on low-dimensional regression problems.
Non-uncertainty based approaches for filtering out poor generations include the realism [Kynkäänniemi et al., 2019], rarity [Han et al., 2023], and anomaly scores [Hwang et al., 2024]. Our work is the first to establish a connection between these scores and uncertainty-based methods, which we hope will inspire the development of even better sample-level metrics in the future. Additionally, a large body of work focuses on specially designed sample-quality scoring models [Gu et al., 2020, Zhao et al., 2024] or, alternatively, on leveraging large pretrained vision-language models (VLMs) [Zhang et al., 2025] for scoring generated images. However, these approaches require either access to sample-quality labels or rely on (expensive) external VLMs. In contrast, our uncertainty-based method requires neither, making it a more accessible and scalable alternative.
[…]
@@ 6 Limitations
While we have demonstrated in Section 4 that semantic likelihood is essential for addressing the over-sensitivity of prior work to background pixels [Kou et al., 2024], our reliance on a pretrained image encoder like CLIP [Radford et al., 2021] limits the applicability of our diffusion uncertainty framework to natural images. Removing the dependence on such encoders would unlock the application our Bayesian framework to other modalities where diffusion models are used, such as molecules [Hoogeboom et al., 2022, Cornet et al., 2024] or text [Gong et al., 2023, Yi et al., 2024]. Exploring whether insights from the literature on uncovering semantic features in diffusion models [Kwon et al., 2023, Luo et al., 2024, Namekata et al., 2024] could help achieve this represents a promising direction for future work.
@@ 7 Conclusion
We introduced generative uncertainty and demonstrated how to estimate it in modern generative models such as diffusion. Our experiments showed the effectiveness of generative uncertainty in filtering out low-quality samples. For future work, it would be interesting to explore broader applications of Bayesian principles in generative modeling beyond detecting poor-quality generations. Promising directions include guiding synthetic data generation and detecting memorized samples. It would also be worthwhile to further investigate the connection between uncertainty-based filtering and classifier(-free) guidance [Ho and Salimans, 2022], as both exhibit similar precision-recall trade-offs.
@@ Acknowledgements.
We thank our reviewers for their thoughtful feedback. This project was generously supported by the Bosch Center for Artificial Intelligence. E.N. did not utilize resources from Johns Hopkins University for this project. S.M. acknowledges funding from the National Science Foundation (NSF) through an NSF CAREER Award IIS-2047418, IIS-2007719, the NSF LEAP Center, the IARPA WRIVA program, and the Hasso Plattner Research Center at UCI. See Appendix for additional disclaimers.
@@
  • In Appendix A, we provide additional figures.
  • In Appendix B.1, we provide more details on our approximation of generative uncertainty (Eq. 3.3).
  • In Appendix C.1, we qualitatively compare our method with BayesDiff [Kou et al., 2024].
  • In Appendix C.2, we perform qualitative ablations on our semantic likelihood (Section 3.3).
  • In Appendix C.3, we demonstrate how to use our generative uncertainty for pixel-wise uncertainty.
@@
  • In Appendix C.4, we show that diffusion’s own likelihood is not useful for filtering out poor samples.
@@
  • In Appendix C.5, we compare our generative uncertainty to realism [Kynkäänniemi et al., 2019] and rarity [Han et al., 2023] scores.
@@
  • In Appendix C.6, we investigate the drop in sample diversity by looking at the average generative uncertainty per conditioning class.
@@
  • In Appendix C.7, we apply our generative uncertainty to detect low-quality samples in a latent flow matching model [Dao et al., 2023].
  • In Appendix D, we provide implementation and experimental details.
Appendix A Additional Figures
Figure 7: Image generation results for $n\in\{6\textrm{K},7\textrm{K},\ldots,11\textrm{K}\}$ filtered samples (out of 12K) for ADM diffusion model [Dhariwal and Nichol, 2021]. Our generative uncertainty outperforms previously proposed uncertainty-based approaches (AU [De Vita and Belagiannis, 2025], BayesDiff [Kou et al., 2024]) in terms of image quality, as indicated by higher FID (left) and precision (middle) scores. We report mean values along with standard deviations over 5 runs with different random seeds.
Appendix B Derivations
B.1 Generative Uncertainty Approximation
To ensure maximal clarity of exposition, we first derive a tractable estimator of $u(\boldsymbol{z})$ (Eq. 4) based on ‘pixel-space’ likelihood (Eq. 6):
[…]
$\mathcal{N}\big{(}\mathbf{x}|\bar{\boldsymbol{g}},\textrm{Diag}\big{(}\frac{1}{M}\sum_{m=1}^{M}g_{\theta_{m}}(\boldsymbol{z})^{2}-\bar{\boldsymbol{g}}^{2}\big{)}+\sigma^{2}I\big{)}=:q_{\textrm{pixel}}(\boldsymbol{z})$
where $\theta_{m}\sim q(\theta|\mathcal{D})$ (Eq. 5) and $\bar{\boldsymbol{g}}=\frac{1}{M}\sum_{m=1}^{M}g_{\theta_{m}}(\boldsymbol{z})$. Note that in step $(1)$ we make use of the usual Monte Carlo (MC) approximation of the Bayesian posterior predictive (Eq. 2) using $M$ samples and in step $(2)$ we approximate a mixture of Gaussian with a single Gaussian using moment-matching. Moreover, we consider only the diagonal of the resulting covariance which we do for efficiency reasons (e.g., for ImageNet 256x256 the full covariance has $\sim 4\cdot 10^{10}$ parameters rendering the diagonal approximation necessary).
A pixel-space generative uncertainty is then obtained by taking the entropy of the resulting Gaussian distribution: $u(\boldsymbol{z})\approx H(q_{\textrm{pixel}}(\boldsymbol{z}))$. However, as we show qualitatively in Appendix C.2, uncertainty based on pixel-space likelihood is not particularly informative about the visual quality of the samples as it is overly sensitive to the background pixels—images with simple backgrounds exhibit low uncertainty, whereas images with ‘cluttered’ background exhibit high uncertainty. This motivates our use of ‘semantic-likelihood’ (Eq. 7) to arrive at the Eq. 3.3:
[…]
$\mathcal{N}\Big{(}\boldsymbol{e}(\mathbf{x})\Big{|}\bar{\boldsymbol{e}},\textrm{Diag}\Big{(}\frac{1}{M}\sum_{m=1}^{M}\boldsymbol{e}_{m}^{2}-\bar{\boldsymbol{e}}^{2}\Big{)}+\sigma^{2}I\Big{)}=:\,q_{\textrm{semantic}}(\boldsymbol{z})$
where $\bar{\boldsymbol{e}}=\frac{1}{M}\sum_{m=1}^{M}\boldsymbol{e}_{m},\;\boldsymbol{e}_{m}=c_{\phi}(g_{\theta_{m}}(\boldsymbol{z})),\;\theta_{m}\sim q(\theta|\mathcal{D})$ and $c_{\phi}$ is a pre-trained feature extractor of choice (e.g., CLIP). With step $(1)$ we again denote the MC approximation (Eq. 2) and step $(2)$ denotes moment-matching (with a diagonal covariance approximation). While the resulting posterior predictive based on the semantic likelihood can not be used to generate samples (since the likelihood is over CLIP features $\mathbf{e}(\mathbf{x})\in\mathcal{S}$ and not data $\mathbf{x}\in\mathcal{X}$), we can still compute its entropy which we use as our final estimate of generative uncertainty $u(\boldsymbol{z})\approx H(q_{\textrm{semantic}}(\boldsymbol{z}))$. As described in Section 3.3 and in Algorithm 1, we first generate a sample using a pretrained model $g_{\hat{\theta}}$ and then use the semantic posterior predictive $q_{\textrm{semantic}}$ solely for uncertainty estimation.
At this point it is important to acknowledge that by changing the likelihood to the semantic one we depart from the traditional Bayesian framework where the same likelihood is used both for finding the posterior $q(\theta|\mathcal{D})$ as well as in the approximation of the posterior predictive. However, we would like to emphasize that image generation using modern diffusion models poses specific challenges, which to the best of our knowledge, have not been addressed within the Bayesian framework yet. One such challenge is due to (extremely) high-dimensional sample spaces. For example, in the case of ImageNet 256x256 the dimensionality is $\sim 2\cdot 10^{5}$. Our use of feature extractor $c_{\phi}$ via the semantic likelihood reduces the dimensionality (down to $512$), potentially making the MC approximation using few samples ($M$) ‘easier’. Another challenge is that using a larger number of MC samples is computational prohibitive, since every additional sample corresponds to generating a new sample with a diffusion model which is costly.
We hope that our promising experimental results based on the semantic likelihood (using a few MC samples only) will encourage the Bayesian community to further investigate the choice of the suitable likelihood in high-dimensional spaces (such as those of natural images) and fill-in the potentially missing theoretical gaps (e.g., due to changing the likelihood ‘post-hoc’).
Appendix C Additional Results
C.1 Qualitative Comparison with BayesDiff